L=100mm
R=100mm
I=100A
1
Expert's answer
2020-02-27T10:09:46-0500

Let's clarify the condition. So, we have


The current "I=100A" flows along the solenoid of length "L=100 mm" and radius "R=100 mm" . Find the induction of the magnetic field in the center of the solenoid. The number of solenoid turns "N=100" and "\mu =1".


"B=\frac {\mu \mu_0IN}{2L}(\cos \alpha_2-\cos \alpha_1)"


"\cos \alpha_1=\cos(90°+45°)=-\sin 45°=-\frac{R}{\sqrt{R^2+R^2}}=-\frac{1}{\sqrt{2}}"


"\cos \alpha_2=\cos(45°)=\frac{1}{\sqrt{2}}"


"B=\frac {\mu \mu_0IN}{2L}(\frac{1}{\sqrt{2}}+\frac{1}{\sqrt{2}})="


"=\frac {1\cdot 4\cdot 3.14 \cdot 10^{-7}\cdot 100\cdot 100}{2\cdot 100}(\frac{1}{\sqrt{2}}+\frac{1}{\sqrt{2}})\approx9\cdot10^{-5} T"